OpenStudy (anonymous):

Wendy takes a multiple-choice quiz that has 20 questions. There are 4 answer choices for each question. What is the probability that she will get at least 2 answers correct by guessing? Round your answer to the nearest hundredth.

OpenStudy (gw2011):

The probability of getting one answer correct by guessing is 1/4. Therefore, the probability of getting 2 answers correct by guessing is: (1/4)(1/4)=1/16=0.0625=6.25%